If a bus is goin 60 kmph and it is delayed 3 hours but if it goes 90 kmph it is one hour early. How long is the route?
Answer by mananth(16946)


If it reaches on time let the number of hours be x
Distance = rate * time

(x+3)*60=(x-1)*90
60x+180=90x-90
-30x=-270
x=-270/-30
x= 9 hours
take any one of the speeds
60 kmph it takes 3 hours more so it travels 12 hours
12 *60 = 720 km
90kmh it takes 1 hour less so itakes 8 hours
8*90 =720 kms
cheers
Ananth